Michelle receives a 10 percent raise every year. Her salary after four such raises
has gone up by what percent?
(A) between 30 and 35 percent (B) between 35 and 40 percent
(C) between 40 and 45 percent (D) between 45 and 50 percent
(E) none of these

Answer by stanbon(75887)

Michelle receives a 10 percent raise every year.
Her salary after four such raises has gone up by what percent?
----
If she starts at salary = x
1st raise she has x + 0.1x = (1.1)x
2nd raise she has (1.1)x + 0.1(1.1x) = (1.1)x(1.1) = (1.1)^2 x
3rd raise she has (1.1)^3 x
4th raise she has (1.1)^4x
----
1.1^4 = 1.4641
1.4641x = x + 0.4641x
% increase = 46.41%
===========================
Cheers,
Stan H.

(A) between 30 and 35 percent (B) between 35 and 40 percent
(C) between 40 and 45 percent (D) between 45 and 50 percent
(E) none of these